OpenStudy (anonymous):

Determine the value of tan 120 (degrees) by using the function of 60 (degrees)

OpenStudy (anonymous):

tan (120) = - tan (60) = - sqrt (3) You must know that tan (60) = sqrt (3)

OpenStudy (anonymous):

Thank you!